# Giant Hall Waves Launched by Superconducting Phase Transition in Pulsars

### v1

@@ 1 Introduction
Early numerical modelling of magnetic field evolution in neutron star interiors neglected the core entirely, and considered the evolution of crust-confined magnetic fields due to Hall drift and Ohmic diffusion (Hollerbach & Rüdiger, 2004; Pons et al., 2009; Viganò et al., 2012; Gourgouliatos & Cumming, 2014). Other studies attempted to model the magnetic field evolution in the core, however the problem is complicated by the effects of proton superconductivity (Ruderman et al., 1998; Ruderman, 2004; Elfritz et al., 2016; Bransgrove et al., 2018).
Nuclear superfluidity in neutron star interiors was predicted by Migdal (1959), eight years before the discovery of the first pulsar. It is now generally accepted that below a critical temperature $T_{c}\sim 10^{9}$ K, conditions in the outer core are suitable for the Cooper pairing of protons and the formation of a type-II superconductor with critical field $H_{\text{c1}}\sim 10^{15}$ G (Baym et al., 1969). An important feature of type-II superconductors is the quantization of the magnetic field into discrete superconducting flux tubes, each of which possess a quantum of flux $\phi_{0}=hc/(2e)$. The bunching of magnetic field lines into discrete flux tubes dramatically modifies the macroscopically averaged Maxwell stress. In particular, the anisotropic components of the Maxwell tensor are larger than $B^{2}/(4\pi)$ by a factor $H_{\text{c1}}/B\sim 10^{3}$ for a typical radio pulsar (Jones, 1975; Easson & Pethick, 1979).
@@ 4.3 Core
Neutron star cores contain degenerate nuclear matter which consists of neutrons, protons and electrons. Protons are expected to form Cooper pairs through the attractive short range component of the strong nuclear force. The properties of the superconductor are determined by its penetration depth $\lambda_{p}=(m_{p}c^{2}/4\pi n_{p}e^{2})^{1/2}$, and coherence length $\xi_{p}\approx(2/\pi k_{F})(\epsilon_{p}/\Delta_{p})$ where $k_{F}$ is the proton Fermi wavevector, $\epsilon_{p}$ the proton Fermi energy, and $\Delta_{p}$ the superconducting energy gap (Baym et al., 1969). Typical densities in the outer core imply $\lambda_{p}/\xi_{p}\gg 1$, so the superconductor is of type-II. Thus, at the time of the phase transition the magnetic field is quantized into an array of discrete superconducting flux tubes which each contain the quantum of flux $\phi_{0}=hc/(2e)$ within a characteristic radius $\lambda_{p}$. The local magnetic field at the center of a flux tube is $\boldsymbol{H}\approx H_{\text{c1}}\boldsymbol{B}/|\boldsymbol{B}|$, with $H_{\text{c1}}\approx\phi_{0}/(\pi\lambda_{p}^{2})\sim 10^{15}$ G, while the field is exponentially suppressed in the volume between flux tubes (Easson & Pethick, 1977). In the neutron star core, we use $\boldsymbol{B}$ to represent the magnetic field which is spatially averaged over many flux tubes such that $|\boldsymbol{B}|=n_{\Phi}\phi_{0}$, where $n_{\Phi}$ is the local area density of flux tubes.
@@ 5 Sample Models
Figure 2: Magnetic field evolution of Model D2. Top row: Snapshots of the magnetic field configuration at $t=0$ kyr, $t=20$ kyr, and $t=200$ kyr. Green curves show poloidal magnetic field lines, and color shows $B_{\phi}$ in units of $10^{14}$ G. Bottom row: Snapshots of the von Mises strain $|\epsilon|=\sqrt{\frac{1}{2}\epsilon_{ij}\epsilon_{ij}}$ at the same times. The crust-core interface and the surface are indicated by the inner and outer dashed white curves, respectively. The axis show distance in units of $10^{6}$ cm.
We present five sample models, which cover a range of initial magnetic fields and different models of superconducting flux tube drift in the core. The parameters of the models are summarised in Table 1. The dynamics of the magnetic field is fascinating, and we provide movies that illustrate the field evolution in all of these models. In addition, we focus on the following quantities: (1) The von Mises strain of the solid crust $|\epsilon|=\sqrt{\frac{1}{2}\epsilon_{ij}\epsilon_{ij}}$ and its maximum value at a given time $|\epsilon|_{\text{max}}$. The latter diagnoses the possibility of star quakes, and is plotted as a function of time in Fig. 5, (2) The braking index $n=\nu\ddot{\nu}/\dot{\nu}^{2}$, where $\nu$, $\dot{\nu}$, and $\ddot{\nu}$ are the pulsar rotation frequency and it’s time derivatives. The braking index is commonly measured by pulsar observers, and is plotted in Fig. 5.
@@
Ohmic decay of electric currents in the closed magnetosphere around the neutron star is fast compared to the internal magnetic evolution studied in this paper. Therefore, outside the neutron star surface, $r>R_{\star}$, we assume a vacuum magnetosphere so that the magnetic field at $R_{\star}$ matches onto a potential field (×$\nabla\times\boldsymbol{B}=0$) at each time step (Bransgrove et al., 2018). We assume that the superconducting phase transition happens at the beginning of the simulation ($t=0$). Therefore, at the crust-core interface we enforce continuity of stress at each time step, as discussed in Section 3, $B_{r}(r_{c}^{+})=B_{r}(r_{c}^{-})$, $B_{\theta}(r_{c}^{+})=H_{\theta}(r_{c}^{-})$, and $B_{\phi}(r_{c}^{+})=H_{\phi}(r_{c}^{-})$, where $r_{c}$ is the radius of the crust-core interface. All of the initial magnetic fields satisfy ×$\nabla\times\boldsymbol{B}=0$ in the crust, so that the resulting evolution is entirely due to the crust-core coupling.
Figure 3: Same as Fig. 2, but for Model M1. In the core, color shows $H_{\phi}$ in units of $10^{14}$ G.
@@ 5.2 Moderate flux tube relaxation (finite $\tau$)
In Model M1 the flux tubes drift due to their own self-tension force (Eq. 18). Flux tubes minimize their energy by increasing their radius of curvature, and shortening their length. Thus, as the flux tubes relax and straighten, poloidal field lines develop strong curvature at the crust-core interface. This produces a powerful current $j_{\phi}$ in the deep crust, which shears field lines in the azimuthally due to the Hall effect. The strong azimuthal shear of fields in the deep crust twists flux tubes, and results in non-zero toroidal fields in the core ($H_{\phi}\neq 0$, Fig 3 middle panel). Currents generated by toroidal fields in the deep crust advect poloidal field lines in the $\hat{\theta}$ direction, pushing together flux tubes of opposite polarity, and forcing reconnection at the crust-core interface at $t\gtrsim 20$ kyr. After $\sim 50$ kyr, all of the flux tube loops have either entered the crust, or reconnected with neighbouring loops (Fig. 3, last panel). The flux tubes which remain in the core extend from north to south hemisphere, and are untwisted ($H_{\phi}=0$, Fig 3 last panel) due to equatorial symmetry. The flux tubes then slowly drift out of the core and enter the crust, in a manner similar to Models D1 and D2.
@@ 6 Discussion
The Hall waves excited by the phase transition also couple to slow changes of the magnetosphere. These changes have an observable effect on the pulsar spin-down (Fig. 5, right), which is often characterized by the braking index $n=\nu\ddot{\nu}/\dot{\nu}^{2}$. Force-free models of the magnetosphere overwhelmingly give $n=3$, regardless of the inclination angle (Goldreich & Julian, 1969; Contopoulos et al., 1999; Gruzinov, 2005; Spitkovsky, 2006). Remarkably, observed pulsars have braking indices of positive and negative sign spanning the range $-10^{8}\lesssim n\lesssim 10^{8}$ (Hobbs et al., 2010; Zhang & Xie, 2012). The implication from both theory and observation is that the pulsar magnetic fields must be strongly evolving on a timescale much shorter than the pulsar’s age. Hall drift in the crust naturally produces anomalous braking indices $n\neq 3$ – this effect was demonstrated by Pons et al. (2012) and Gourgouliatos & Cumming (2015). The braking indices measured in our simulations ($|n|\lesssim 50$, Fig. 5) are similar to the braking indices of observed pulsars younger than $\sim 200$ kyr which span $|n|\lesssim 100$ (Zhang & Xie, 2012; Parthasarathy et al., 2019, 2020; Lower et al., 2021). Our simulations produce $n$ and $\ddot{\nu}$ of positive and negative sign after the first $\sim 50-100$ kyr, similar to observed pulsars (Hobbs et al., 2010).

### v2

@@ 1 Introduction
Early numerical modelling of magnetic field evolution in neutron star interiors neglected the core entirely, and considered the evolution of crust-confined magnetic fields due to Hall drift and Ohmic diffusion (Hollerbach & Rüdiger, 2004; Pons et al., 2009; Viganò et al., 2012; Gourgouliatos & Cumming, 2014; Marchant et al., 2014). Other studies attempted to model the magnetic field evolution in the core, however the problem is complicated by the effects of proton superconductivity (Ruderman et al., 1998; Ruderman, 2004; Elfritz et al., 2016; Bransgrove et al., 2018).
Nuclear superfluidity in neutron star interiors was predicted by Migdal (1959), eight years before the discovery of the first pulsar. It is now generally accepted that below a critical temperature $T_{c}\sim 10^{9}-10^{10}$ K, conditions in the outer core are suitable for the Cooper pairing of protons and the formation of a type-II superconductor with critical field $H_{\text{c1}}\sim 10^{15}$ G (Baym et al., 1969). An important feature of type-II superconductors is the quantization of the magnetic field into discrete superconducting flux tubes, each of which possess a quantum of flux $\phi_{0}=hc/(2e)$. The bunching of magnetic field lines into discrete flux tubes dramatically modifies the macroscopically averaged Maxwell stress. In particular, the anisotropic components of the Maxwell tensor are larger than $B^{2}/(4\pi)$ by a factor $H_{\text{c1}}/B\sim 10^{3}$ for a typical radio pulsar (Jones, 1975; Easson & Pethick, 1979).
@@ 4.3 Core
Neutron star cores contain degenerate nuclear matter which consists of neutrons, protons and electrons. Protons are expected to form Cooper pairs through the attractive short range component of the strong nuclear force. The properties of the superconductor are determined by its penetration depth $\lambda_{p}=(m_{p}c^{2}/4\pi n_{p}e^{2})^{1/2}$, and coherence length $\xi_{p}\approx(2/\pi k_{F})(\epsilon_{p}/\Delta_{p})$ where $k_{F}$ is the proton Fermi wavevector, $\epsilon_{p}$ the proton Fermi energy, and $\Delta_{p}$ the superconducting energy gap (Baym et al., 1969). Typical densities in the outer core imply $\lambda_{p}/\xi_{p}\gg 1$, so the superconductor is of type-II. Thus, at the time of the phase transition the magnetic field is quantized into an array of discrete superconducting flux tubes which each contain the quantum of flux $\phi_{0}=hc/(2e)$ within a characteristic radius $\lambda_{p}$. The local magnetic field inside a flux tube is $\boldsymbol{H}\approx H_{\text{c1}}\boldsymbol{B}/|\boldsymbol{B}|$, with $H_{\text{c1}}\approx\phi_{0}/(\pi\lambda_{p}^{2})\sim 10^{15}$ G, while the field is exponentially suppressed in the volume between flux tubes (Easson & Pethick, 1977). In the neutron star core, we use $\boldsymbol{B}$ to represent the magnetic field which is spatially averaged over many flux tubes such that $|\boldsymbol{B}|=n_{\Phi}\phi_{0}$, where $n_{\Phi}$ is the local area density of flux tubes.
@@ 5 Sample Models
Figure 2: Magnetic field evolution of Model D2. Top row: Snapshots of the magnetic field configuration at $t=0$ kyr, $t=20$ kyr, and $t=200$ kyr. Green curves show poloidal magnetic field lines, and color shows $B_{\phi}$ in units of $10^{14}$ G. Bottom row: Snapshots of the von Mises strain $|\epsilon|=\sqrt{\frac{1}{2}\epsilon_{ij}\epsilon_{ij}}$ at the same times. The crust-core interface and the surface are indicated by the inner and outer dashed white curves, respectively. The axis show distance in units of $10^{6}$ cm. An animated version of this figure, showing the magnetic field evolution from $t=0$ kyr to $200$ kyr, is available in the online article.
We present five sample models, which cover a range of initial magnetic fields and different models of superconducting flux tube drift in the core. The parameters of the models are summarised in Table 1. The dynamics of the magnetic field is fascinating, and we provide movies that illustrate the field evolution corresponding to Figures 2, 3, and 4. In addition, we focus on the following quantities: (1) The von Mises strain of the solid crust $|\epsilon|=\sqrt{\frac{1}{2}\epsilon_{ij}\epsilon_{ij}}$ and its maximum value at a given time $|\epsilon|_{\text{max}}$. The latter diagnoses the possibility of star quakes, and is plotted as a function of time in Fig. 5, (2) The braking index $n=\nu\ddot{\nu}/\dot{\nu}^{2}$, where $\nu$, $\dot{\nu}$, and $\ddot{\nu}$ are the pulsar rotation frequency and it’s time derivatives. The braking index is commonly measured by pulsar observers, and is plotted in Fig. 5.
@@
Ohmic decay of electric currents in the closed magnetosphere around the neutron star is fast compared to the internal magnetic evolution studied in this paper. Therefore, outside the neutron star surface, $r>R_{\star}$, we assume a vacuum magnetosphere so that the magnetic field at $R_{\star}$ matches onto a potential field (×$\nabla\times\boldsymbol{B}=0$) at each time step (Bransgrove et al., 2018). We assume that the phase transition happens at $t=0$, so that the entire core is superconducting throughout the simulation. Therefore, at each time step we enforce the continuity of stress at the crust-core interface, as discussed in Section 3, $B_{r}(r_{c}^{+})=B_{r}(r_{c}^{-})$, $B_{\theta}(r_{c}^{+})=H_{\theta}(r_{c}^{-})$, and $B_{\phi}(r_{c}^{+})=H_{\phi}(r_{c}^{-})$, where $r_{c}$ is the radius of the crust-core interface. All of the initial magnetic fields satisfy ×$\nabla\times\boldsymbol{B}=0$ in the crust, so that the resulting evolution is entirely due to the crust-core coupling. We choose initial magnetic fields with $B\ll H$ similar to real pulsars, so that the magnetic stress in the core is much larger than the magnetic stress in the crust. This ensures that the Hall waves have a very large amplitude. For magnetars with $B\sim H$, the magnetic stress in the crust and the core is similar, and the effect of the phase transition is much less dramatic.
Figure 3: Same as Fig. 2, but for Model M1. In the core, color shows $H_{\phi}$ in units of $10^{14}$ G. An animated version of this figure, showing the magnetic field evolution from $t=0$ kyr to $200$ kyr, is available in the online article.
@@ 5.1 Infinitely fast flux tube relaxation ($\tau=0$)
Figure 4: Same as Fig. 2, but for Model M3. An animated version of this figure, showing the magnetic field evolution from $t=0$ kyr to $200$ kyr, is available in the online article.
@@ 5.2 Moderate flux tube relaxation (finite $\tau$)
In Model M1 the flux tubes drift due to their own self-tension force (Eq. 18). Flux tubes minimize their energy by increasing their radius of curvature, and shortening their length. Thus, as the flux tubes relax and straighten, poloidal field lines develop strong curvature at the crust-core interface. This produces a powerful current $j_{\phi}$ in the deep crust, which shears field lines in the azimuthal direction due to the Hall effect. The strong azimuthal shear of fields in the deep crust twists flux tubes, and results in non-zero toroidal fields in the core ($H_{\phi}\neq 0$, Fig 3 middle panel). Currents generated by toroidal fields in the deep crust advect poloidal field lines in the $\hat{\theta}$ direction, pushing together flux tubes of opposite polarity, and forcing reconnection at the crust-core interface at $t\gtrsim 20$ kyr. After $\sim 50$ kyr, all of the flux tube loops have either entered the crust, or reconnected with neighbouring loops (Fig. 3, last panel). The flux tubes which remain in the core extend from north to south hemisphere, and are untwisted ($H_{\phi}=0$, Fig 3 last panel) due to equatorial symmetry. The flux tubes then slowly drift out of the core and enter the crust, in a manner similar to Models D1 and D2.
@@ 6 Discussion
The Hall waves excited by the phase transition also couple to slow changes of the magnetosphere. These changes have an observable effect on the pulsar spin-down (Fig. 5, right), which is often characterized by the braking index $n=\nu\ddot{\nu}/\dot{\nu}^{2}$. Force-free models of the magnetosphere overwhelmingly give $n=3$, regardless of the inclination angle (Goldreich & Julian, 1969; Contopoulos et al., 1999; Gruzinov, 2005; Spitkovsky, 2006). Remarkably, observed pulsars have braking indices of positive and negative sign spanning the range $-10^{8}\lesssim n\lesssim 10^{8}$ (Hobbs et al., 2010; Zhang & Xie, 2012). The implication from both theory and observation is that the pulsar magnetic fields must be strongly evolving on a timescale much shorter than the pulsar’s age. Hall drift in the crust naturally produces anomalous braking indices $n\neq 3$ because the dipole moment becomes time-dependent – this effect was demonstrated by Pons et al. (2012) and Gourgouliatos & Cumming (2015). The braking indices measured in our simulations ($|n|\lesssim 50$, Fig. 5) are similar to the braking indices of observed pulsars younger than $\sim 200$ kyr which span $|n|\lesssim 100$ (Zhang & Xie, 2012; Parthasarathy et al., 2019, 2020; Lower et al., 2021). Our simulations produce $n$ and $\ddot{\nu}$ of positive and negative sign after the first $\sim 50-100$ kyr, similar to observed pulsars (Hobbs et al., 2010).